OpenStudy (anonymous):

The slope-intercept equation of a line is y = -7x - 2. What are the slope and y-intercept of the line? A.The slope is -7 and the y-intercept is (0, -2). B.The slope is -7 and the y-intercept is (0, 2). C.The slope is -2 and the y-intercept is (0, 7). D.The slope is -2 and the y-intercept is (0, -7).

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thx

OpenStudy (usukidoll):

hmm? we have the equation in the form of y=mx+b where m is the slope and b is the y-intercept

OpenStudy (usukidoll):

so if we're given y =-7x-2 and it's the slope intercept form of y=mx+b. What is m and what is b?

OpenStudy (usukidoll):

I will explain to you why it's NOT C If option C were to hold true, we would've gotten the equation y = -2x+7 The original equation, y =-7x-2 contains 2 negative numbers which is -7 and -2 Option C contains 1 positive number and 1 negative number which is -2 and 7

OpenStudy (anonymous):

wait wht

OpenStudy (usukidoll):

try this again. if we're given y =-7x-2 and it's the slope intercept form of y=mx+b. What is m and what is b?

OpenStudy (anonymous):

m is -7 and b is 2

OpenStudy (anonymous):

right?

OpenStudy (usukidoll):

the m is correct.. the b isn't because it's a negative number

OpenStudy (anonymous):

so -2

OpenStudy (usukidoll):

yes so m = -7 and b =-2 It makes sense because we were given y =-7x-2

OpenStudy (usukidoll):

m = slope and b = y-intercept therefore slope = -7 and y-intercept = -2

OpenStudy (anonymous):

so a

OpenStudy (usukidoll):

"A.The slope is -7 and the y-intercept is (0, -2)." correct

OpenStudy (anonymous):

thx can you help with one more

OpenStudy (usukidoll):

Be careful for those direct answer guys.. they are just doing it to gain medals and have no idea what they are doing. That previous user was lucky that it was the wrong answer. Had it been right, that user would've violated the Code of Conduct which is to never give direct answers and to guide users through the learning process. yeah sure.

OpenStudy (anonymous):

If two lines are parallel and one has a slope of 8, what is the slope of the other line?

OpenStudy (anonymous):

would it be the same

OpenStudy (anonymous):

A.8 B.-1/8 C.-8 D.1/8

OpenStudy (usukidoll):

parallel lines look like this  they don't intersect .

OpenStudy (anonymous):

ik so the slope would be 8

OpenStudy (usukidoll):

the slope is positive though. we are given that there are two parallel lines, but one of them has the slope of 8

OpenStudy (anonymous):

its either a or d

OpenStudy (anonymous):

wait its -8 :P

OpenStudy (usukidoll):

really? can you tell me why it's -8? I'm curious ^^

OpenStudy (usukidoll):

the problem is ... we are not given perpendicular lines.. we have 2 parallel lines

OpenStudy (anonymous):

so its not -8

OpenStudy (usukidoll):

for perpendicular lines we can flip the slope over. perpendicular slopes have opposite signs.

OpenStudy (anonymous):

soo for parallel the slope is the same

OpenStudy (usukidoll):

parallel means that the slopes are the same parallel lines are just lines that don't intersect. If I had 2 perpendicular lines I will be intersecting at some point

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

??

OpenStudy (usukidoll):

sorry I was responding to another question . yes it will be the same . since the slope is 8 we have 8 :D

OpenStudy (anonymous):

thx